OpenStudy (anonymous):

How do you verify the trig identity : cos squared theta - sin squared theta =2 cos squared theta - 1 ?

OpenStudy (solomonzelman):

I will change to using x's that is more convenient for me. \(\large\color{blue}{ \cos^2x -\sin^2x=2\cos^2x-1 }\) use the fact \(\large\color{black}{ \sin^2x=1-\cos^2x }\)

OpenStudy (anonymous):

Thank you,I understand now